OpenStudy (prince1342):

Given the functions f(x) = 4x2 − 1, g(x) = x2 − 8x + 5, and h(x) = –3x2 − 12x + 1, rank them from least to greatest based on their axis of symmetry.

OpenStudy (prince1342):

g(x), h(x), f(x) f(x), h(x), g(x) g(x), f(x), h(x) h(x), f(x), g(x)

sam (.sam.):

Use \(x=-\frac{b}{2a}\) to find the symmetries, using \(ax^2+bx+c\) as reference, For f(x), since b=0, x=0. For g(x), \(x=-\frac{(-8)}{2(1)}=4\) For h(x). \(x=-\frac{(-12)}{2(-3)}=-2\) Rank them now

OpenStudy (prince1342):

Making it D?

sam (.sam.):

Yes

OpenStudy (prince1342):

can you help with another?

sam (.sam.):

Just post them